OpenStudy (anonymous):

Which expressions show how 8 • 54 can be rewritten using the distributive property? Choose all answers that are correct. A. 8 • 50 + 4 B. 8 • 50 + 8 • 4 C. 8 • 60 – 8 • 6 D. 8 • 60 + 8 • 4 I know one of them is B

OpenStudy (anonymous):

A is incorrect since 8 • 50 + 4=404, and not 432

OpenStudy (anonymous):

C is correct

OpenStudy (anonymous):

B is correct since 8 • 50 + 8 • 4= 432

OpenStudy (anonymous):

C is correct since 8 • 60 – 8 • 6=432

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

D is incorrect since 8 • 60 + 8 • 4=512

OpenStudy (anonymous):

ighht thank ya :D